You have twice as many 20 cent coins as 10 cents coins and half the number of 5 cent coins as 10 cent coins. If you have R4,20 altogether, find out how many of each coin denomination do you have?


1
Expert's answer
2021-10-08T03:59:50-0400

Let the number of 10 cent coins =x, 20 cent coins =2x, and 5 cent coins = 1/2x

"10x+20(2x)+5(1\/2x)=420"

"52.5x=420"

"x=8"

20 cent coins are "2\times 8=16"

10 cent coins are "8\times 1=8"

5 cent coins are "1/2\times 8=4"